6 g/cm^3 is how much in lbs/ ft^3
Answer by ptaylor(1894)

6 g/cm^3 is how much in lbs/ ft^3

1 gram = 0.002 lbs
6gm=0.012 lb
1 centimeter = 0.033 feet
1cm^3=(0.033)^3=0.00004 ft^3
So, 6gm/cm^3=0.012lb/0.00004ft^3= (0.012/0.00004) lb/ft^3=300lb/ft^3----WoW!!!!
CHECK MY MATH

Hope this helps---ptaylor